# Flexible Music-Conditioned Dance Generation with Style Description Prompts

### v1

@@ Abstract
Dance plays an important role as an artistic form and expression in human culture, yet the creation of dance remains a challenging task. Most dance generation methods primarily rely solely on music, seldom taking into consideration intrinsic attributes such as music style or genre. In this work, we introduce Flexible Dance Generation with Style Description Prompts (DGSDP), a diffusion-based framework suitable for diversified tasks of dance generation by fully leveraging the semantics of music style. The core component of this framework is Music-Conditioned Style-Aware Diffusion (MCSAD), which comprises a Transformer-based network and a music Style Modulation module. The MCSAD seemly integrates music conditions and style description prompts into the dance generation framework, ensuring that generated dances are consistent with the music content and style. To facilitate flexible dance generation and accommodate different tasks, a spatial-temporal masking strategy is effectively applied in the backward diffusion process. The proposed framework successfully generates realistic dance sequences that are accurately aligned with music for a variety of tasks such as long-term generation, dance in-betweening, dance inpainting, and etc. We hope that this work has the potential to inspire dance generation and creation, with promising applications in entertainment, art, and education.
@@
Pedestrian detection, hard example mining.
@@ I Introduction
Music has always been a pivotal force in driving dance movement, whether through the rhythmic beats of traditional folk dances or the innovative choreography of contemporary performances. With the continuous development of artificial intelligence technology, there has been a growing interest in exploring the intersection between music and dance. Automated dance generation has garnered immense interest, offering fresh perspectives on the intricate relationship between music and dance, while simultaneously inspiring innovative dance creation and synthesis.
Music-conditioned dance generation aims to automatically produce dance movements that are in harmony with musical inputs. It is closely intertwined with human motion synthesis [1], which involves creating realistic and natural human motions. The main distinction lies in the fact that dance generation, which necessitates a profound analysis of music’s characteristics, is a cross-modal generation task. This task is also a typical problem in generating human motions based on conditional signals [2], which is extremely challenging due to the implicit relationship between human motion and conditional signals.
Figure 1: The illustration of dance generation with style description prompts.
Recently, advanced generative models such as Generative Adversarial Networks (GANs) [3], Variational Autoencoders (VAEs) [4], Cross-modal Transformer [5] and Diffusion [6] are employed to address music-conditioned dance generation. Although these models are capable of successfully generating high-quality and diverse dances from music signals, they overlook the significance of inherent music attributes, such as style or genre, in the dance generation process. Indeed, choreographers consider the music style to create a dance that is aesthetically pleasing to human beings. Music styles set the tone for the creative direction and style of dance generation. Different genres of music, such as classical and pop, offer varied dance forms. Therefore, careful incorporating music style is crucial in music-conditioned dance generation.
@@
Another important aspect that cannot be neglected in dance generation is the flexibility. Dance, as a dynamic art form, often requires meticulous adjustments and refinements throughout its creation and performance phases. The ability to flexibly edit dances can not only enhance artistic expression but also enable dancers to respond effectively to user feedback. A practical dance generation approach ought to possess sufficient flexibility to cater to a diverse range of tasks.
@@
To address above issues, we introduce a novel Music-Conditioned Dance Generation with Style Description Prompts (DGSDP), as illustrated in Figure 2. The DGSDP mainly comprises Music-Conditioned Style-Aware Diffusion (MCSAD) and Style Modulation with Description Prompts. The MCSAD is a transformer-based diffusion framework that seemly integrates music conditions and style description prompts into the dance generation. To fully utilize the style information for dance generation, we introduce a Style Modulation module and explore three types of style embeddings: one-hot encoding, genre embedding, and style description prompts. The proposed diffusion model is jointly trained with classifier-free guidance. To accommodate different dance generation tasks, a spatial-temporal masking is integrated in the backward diffusion process. To better align practical applications and facilitate future research, we design a range of experimental settings and establish benchmarks of different approaches under these settings.
The main contributions are as follows: (1) We propose flexible music-conditioned Dance Generation with Style Description Prompts (DGSDP) which deliberately incorporates style description prompts to enhance dance generation with styles. (2) We introduce Music-Conditioned Style-Aware Diffusion (MCSAD) which primarily comprises a Transformer-based dance generation network and a novel style modulation module. (3) To accommodate diverse dance generation in practical scenarios, we design a flexible dance generation algorithm by leveraging a spatial-temporal masking strategy. (4) The proposed method exhibits state-of-the-art performance across various dance generation tasks, encompassing long-term dance generation, dance in-betweening and dance inpainting.
@@ II Related Work
II-A Stochastic Human Motion Prediction
Human Motion Prediction (HMP) can be categorized into deterministic prediction and stochastic prediction. Stochastic HMP has become popular with the emergence of generative models such as Variational Auto-Encoders (VAE) and Diffusion. Early work mostly relies on autoregressive models, which have the disadvantage of error accumulation. Yan $et$ $al.$ [7] propose a network that can generate the entire sequence directly by transforming from a series of latent vectors sampled from a Gaussian process (GP). Considering the inherent stochasticity in future motions, Kundu $et$ $al.$ [8] present a new probabilistic generative technique by introducing a random extrinsic factor. Aliakbarian $et$ $al.$ [9] propose a model with the ability to produce multiple probable future pose sequences by forcing the model to consider the noise. Mao $et$ $al.$ [10] develop a unified deep generative network and propose a pose prior and a joint angle constraint for producing smooth pose sequences. Instead of defining a combined loss as in the proposed work, Ma $et$ $al.$ [11] achieve the target of balancing the diversity sampling and the accuracy sampling by presenting a unified multi-objective conditional variational autoencoder. Given action labels and motion history, Mao $et$ $al.$ [12] design a VAE-based model which bridges the gap between stochastic human motion prediction and motion synthesis. Diffusion models have been applied for HMP. Chen $et$ $al.$ [13] present a diffusion-based framework with masked completion. Wei $et$ $al.$ [14] propose a diffusion probabilistic model to achieve diversity by incorporating a new noise at each diffusion timestep.
II-B Music-Conditioned Dance Generation
Generating dances that match the music temporally and aesthetically is challenging, and various models such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Transformers, and Diffusion are employed to tackle this task. Li $et$ $al.$ [5] present a model involving a deep cross-modal transformer block with full-attention. Valle-Pérez $et$ $al.$ [15] introduce an innovative probabilistic autoregressive framework, taking into account prior poses and musical context. Au $et$ $al.$ [16] propose a motion generation framework, which has the capability to choreograph high-quality dance movements over a Dynamic Graph. Huang $et$ $al.$ [17] construct a transformer-based framework to generate dance with a specific genre. Kim $et$ $al.$ [3] propose a transformer-based conditional GAN framework by sampling from each latent representation of different dance genres. To generate plausible long-term dance sequences, Li $et$ $al.$ [18] introduce a two-stage process: a key pose generation and an in-between parametric motion curve prediction. Li $et$ $al.$ [19] propose a architecture with two components: a choreographic memory for encoding and quantizing dance poses and an actor-critic GPT for composing units to dance. Tseng $et$ $al.$ [6] utilize a diffusion model which enables powerful editing capabilities. Alexanderson $et$ $al.$ [20] use diffusion models for audio-driven human motion generation, using Conformers. Recently, Gong $et$ $al.$ [21] incorporate both text and music for dance generation by using a cross-modal transformer to encode multi-modal features. Unlike the above approaches, we concentrate on flexible music-conditioned dance generation and enhance generated results by integrating style information within a diffusion framework.
II-C Text-Conditioned Motion Synthesis
Text is often employed as instructions for generating motions, and similarly, it can be used as an additional condition to guide the generation of dance. Ahuja $et$ $al.$ [22] introduce a model that learns a joint embedding space of pose and natural language. Instead of long natural language, Guo $et$ $al.$ [23] propose a novel VAE framework to iteratively generate human motion sequences given a prescribed action type. Then, action is also used by Petrovich  [24], who address the problem of generating human motion sequences conditional on actions. Models in previous work all focus on an entire body. Ghosh $etal.$ [25] explore a joint mapping between 3D pose sequences and textual descriptions by presenting a hierarchical two-stream sequential model. Guo $et$ $al.$ [26] use a two-stage approach: text2length sampling and text2motion generation to generate plausible motions. Guo $et$ $al.$ [27] propose a motion token representation, which can be used in text2motion and motion2text tasks. Instead of synthesizing deterministic motions, Petrovich $et$ $al.$ [28] use variational autoencoder and a text encoder to produce diverse 3D human motions. Inspired by the powerful generative capabilities of the Diffusion model, Guy $et$ $al.$ [29] introduce a classifier-free diffusion-based generative model, which achieves impressive results in text-to-motion tasks. Zhang $et$ $al.$ [30] investigate a simple conditional generative framework based on VQ-VAE and GPT from textual descriptions. Zhang $et$ $al.$ [31] propose a diffusion model-based framework with multi-level manipulation, which can generate arbitrarily long motions and respond to fine-grained instructions on joints.
III Preliminaries
III-A Human Motion Synthesis
@@
III-B Diffusion-Based Dance Generation
Following the definition of DDPM [33], diffusion can be described as a Markov noise process, where the latent variable ${\{\sigma_{t}\}}^{T}_{t=0}$ follows a forward noise process $q(\sigma_{t}|x)$ and $x\sim p(x)$ is taken from the data distribution. The forward noise process is defined as:
$q(\sigma_{t}|x)=\sqrt{\overline{\alpha}_{t}}x+(1-\overline{\alpha}_{t})\epsilon,$ (4)
where $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$ and $\overline{\alpha}_{t}\in(0,1)$ are constants that are monotonically decreasing.
Given the music condition $c$, the model learns to estimate the ground-truth human motion $\hat{x}_{\theta}(\sigma_{t},t,c)\approx x$ at all time moments using the model parameter $\theta$ through a backward diffusion process. We optimize $\theta$ using the simple objective loss [33]:
${\mathcal{L}_{d}={\mathbb{E}}_{x,t}[{||x-{\hat{x}}_{\theta}(\sigma_{t},t,c)||}^{2}_{2}]}.$ (5)
During each denoising time step $t$, the network $\hat{x}(\cdot)$ predicts the denoised samples and adds noise $\hat{\sigma}_{t-1}\sim q(\hat{x}_{\theta}(\hat{\sigma}_{t},t,c),t-1)$, the process is repeated from $T$ to 00 and terminates when $t=0$. We employs a classifier-free guidance [34] for training, a technique commonly utilized in diffusion-based models [35, 36, 37, 29, 31]. Following the approach [34], we incorporate classifier-free guidance by introducing a low probability (e.g. 20%) of randomly replacing $c=\emptyset$ during training. The guided inference is formulated as a weighted combination of unconditional and conditional generated samples:
$\widetilde{x}(\hat{\sigma}_{t},t,c)=w\cdot\hat{x}(\hat{\sigma}_{t},t,c)+(1-w)\cdot\hat{x}(\hat{\sigma}_{t},t,\emptyset),$ (6)
@@ III-B Diffusion-Based Dance Generation
The overall training loss consists of a weighted combination of the primary simple loss and the supplementary auxiliary loss:
@@
where $\lambda_{j}$, $\lambda_{v}$ and $\lambda_{f}$ are weighted coefficients.
Figure 2: The proposed Dance Generation with Style Description Prompts (DGSDP) pipeline. The model is fed with a noisy motion sequence $X_{T}$ of length $N$ in a noising step $T$, along with conditioning music $c$, style description prompts $s$ and $T$ itself. The music embedding serves as input for the cross-attention module. During the inference process, DGSDP samples noise $X_{T}$ given the conditions $c$ and $s$ to predict the clean sample $\hat{X}$. Subsequently, it diffuses this sample back to $X_{T-1}$, and repeats this iterative process until $t=0$ is reached.
IV Method
We introduce a novel Dance Generation with Style Description Prompts (DGSDP), which is illustrated in Figure 2. The DGSDP mainly includes Style Modulation with Description Prompts and Style Modulation and Music-Conditioned Style-Aware Diffusion. Enhanced by a spatial-temporal masking strategy, it offers flexibility in addressing diverse dance generation tasks and possesses the capability to generate long-term dances of arbitrary lengths.
IV-A Style Description Prompts
Genre Embedding: To obtain a more semantic representation, we utilize CLIP [38] to extract features for the words associated with the music genre, and refer to it as Style Type Embedding. This embedding captures deep contextual relationships within the textual genre, resulting in more accurate feature representations.
Style Description Prompts: Large-scale language models pre-trained on massive text [39] demonstrate impressive capabilities for text generation tasks [40]. To get a more detailed description about characteristics of the music style, we use GPT-3 [41] to generate style description prompts for dance. We choose the following function to acquire the style description prompts: ”Please generate a detailed description of the dance [$g$], including the characteristics of the dance in terms of body movement.”, where $g$ is the dance genre.
IV-B Music-Conditioned Style-Aware Diffusion
We extend the Human Motion Diffusion [33] by incorporating style modulation with description prompts, resulting in Music-Conditioned Style-Aware Diffusion (MCSAD). In addition to the music condition $c$, our model also includes the style description prompts $s$. The training objective is:
where $\hat{x}(\cdot)$ denotes the transformer-based decoder block, which primarily comprises self-attention, feature-wise linear modulation (FiLM) [42], cross-attention, and Style Modulation.
The Style Modulation layer aims to increase the influence of style features on the generated dances. It draws inspiration from the adaptive instance normalization technique utilized in StyleGAN [43, 44]. This module $\mathrm{SM}(\cdot)$ allows for the incorporation of additional textual information into existing transformer models. Formally,
where $z$ refers to the input, $s$ represents the music style description prompts, $r$ is a scaling factor, $\mathrm{FC}(\cdot)$ denotes the fully connected layer, $d_{s}$ is the embedding dimension and $d_{z}$ is the hidden dimension.
With the style description prompts, the guided inference is:
$\widetilde{x}(\hat{\sigma}_{t},t,c,s)=w\cdot\hat{x}(\hat{\sigma}_{t},t,c,s)+(1-w)\cdot\hat{x}(\hat{\sigma}_{t},t,\emptyset,s).$ (10)
The detailed training algorithm is shown in Algorithm 1.
@@ IV-B Music-Conditioned Style-Aware Diffusion
      Add noise on $x$ with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t}$: $\sigma_{t}=\sqrt{\overline{\alpha}_{t}}x+(1-\overline{\alpha}_{t})\epsilon$.
@@
      Calculate the losses $L_{j}$, $L_{v}$, and $L_{f}$ defined in Eq. (1-3).
      Calculate the training loss $L_{d}$ defined in Eq. (8).
@@
Figure 3: Flexible dance generation with spatial-temporal masking. For the known sequence, we add noise on it to obtain noisy sequence at timestep $t-1$ directly. For unknown sequences, we first use the trained network ${\hat{x}}_{\theta}$ to predict the motion at the timestep 0, and then we add noise on it to obtain noisy sequence at timestep $t-1$ The mask is two-dimensional and allows for control in both the temporal dimension (frame) and spatial dimension (joint).
IV-C Flexible Generation with Masking
During inference, the proposed method generates an estimated sequence from the initial noisy sequence $\sigma_{T}\sim\mathcal{N}(0,\boldsymbol{I})$, denoises it to $\hat{\sigma}_{T-1}$, and iterates this process until $t=0$.

### v2

# Controllable Dance Generation with Style-Guided Motion Diffusion

@@ Abstract
Dance plays an important role as an artistic form and expression in human culture, yet automatically generating dance sequences is a significant yet challenging endeavor. Existing approaches often neglect the critical aspect of controllability in dance generation. Additionally, they inadequately model the nuanced impact of music styles, resulting in dances that lack alignment with the expressive characteristics inherent in the conditioned music. To address this gap, we propose Style-Guided Motion Diffusion (SGMD), which integrates the Transformer-based architecture with a Style Modulation module. By incorporating music features with user-provided style prompts, the SGMD ensures that the generated dances not only match the musical content but also reflect the desired stylistic characteristics. To enable flexible control over the generated dances, we introduce a spatial-temporal masking mechanism. As controllable dance generation has not been fully studied, we construct corresponding experimental setups and benchmarks for tasks such as trajectory-based dance generation, dance in-betweening, and dance inpainting. Extensive experiments demonstrate that our approach can generate realistic and stylistically consistent dances, while also empowering users to create dances tailored to diverse artistic and practical needs. Code is available on Github: https://github.com/mucunzhuzhu/DGSDP.
@@
Controllable Human Motion Synthesis, Dance Generation
@@ I Introduction
Dance is a universal form of human expression, combining rhythmic body movements with artistic and cultural elements. The ability to automatically generate realistic and stylistically consistent dance movements has significant potential applications in virtual avatars, video game design and film production. However, dance generation presents unique challenges due to the intricate spatial-temporal dynamics of human motion, the variability of dance styles, and the need for synchronization with external inputs, such as music.
Figure 1: The illustration of controllable dance generation with style prompts and user-provided constraints. We use a diffusion-based model as an example, with controlled constraints serving as guidance for the diffusion process.
Controllable generation is an essential aspect of AI-generated content, enabling users to direct and refine the creative process to meet specific requirements. In image generation, models like ControlNet [1] enable users to manipulate image content through various conditions and prompts. In the realm of motion generation, control mechanisms are crucial for aligning generated content with external constraints [2, 3]. Without controllability, generative models risk producing human motions that lack alignment with user intent or context-specific requirements.
Dance generation, typically conditioned on music, aims to synthesize realistic and expressive dance movements based on input conditions. With the rise of deep learning, data-driven models have significantly advanced this field. Generative Adversarial Networks (GANs) [4, 5] and Transformer-based architectures [6, 7, 8] have been adopted to capture the sequential nature of dance movements. However, these models suffer from mode collapse, where the generator produces limited variations of motions, making them less effective for generating diverse and smooth motions.
Inspired by the success of diffusion-based models in human motion synthesis, several recent works have applied diffusion models to dance generation [9, 10, 11, 12]. Diffusion models naturally avoid mode collapse and can generate diverse and high-quality motions. By iteratively refining motions from noise, these approaches produce smooth and realistic motion sequences based on the music input. However, they overlook the importance of style in generating expressive dances.
@@
Another important aspect that cannot be neglected in dance generation is the controllability. Dance, as a dynamic art form, often requires meticulous adjustments and refinements throughout its creation and performance phases. The ability to flexibly control dances can not only enhance artistic expression but also enable dancers to respond effectively to user feedback. A practical dance generation approach ought to possess sufficient flexibility to cater to a diverse range of tasks. To the best of our knowledge, there are no established benchmarks for controllable dance generation.
Since style encompasses characteristics such emotional tone, the lack of explicit style modeling limits current approaches’ ability to adapt to user-defined stylistic preferences or generate expressive dance movements. Style modeling is crucial for unlocking new possibilities in controllable dance generation. To address this gap, we explicitly integrate style prompts in the diffusion-based framework, as illustrated in Figure 1. We also incorporate user-provided constraints as guidance to control the generated dance movements.
@@
To this end, we introduce Style-Guided Motion Diffusion (SGMD) for controllable dance generation. The SGMD is a transformer-based diffusion framework that seemly integrates music conditions and style prompts into the dance generation. A lightweight Style Modulation module is introduced to enable the integration of style information without affecting the content features. A Spatial-Temporal Masking mechanism is designed to refine dance movements to adhere to specific temporal and joint constraints. To fully utilize the style information, we explore three types of style prompts: one-hot encoding, genre name, and style description prompts. The description prompt is initially generated by GPT-3 through multiple trials and manually refined to meet the requirements. To accommodate different dance generation tasks, a spatial-temporal masking is integrated in the backward diffusion process. To better align practical applications, we setup several settings and establish benchmarks under these settings.
Our contributions can be summarized as follows:
  • We study the important yet underexplored task of controllable dance generation and propose a style-guided motion diffusion to address it.
  • We propose a lightweight style modulation module and a spatial-temporal masking to integrate style prompts and enforce spatial constraints, respectively, within the diffusion-based dance generation model.
  • We set up new benchmarks for controllable dance generation tasks, such as trajectory-based dance generation, dance in-betweening, and dance inpainting, while our approach achieves state-of-the-art performance across a variety of dance generation and editing tasks.
@@ II Related Work
II-A Music-Conditioned Dance Generation
Generating dances synchronized with music in both timing and aesthetics remains a challenging task, with models such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Transformers being applied in this domain. A GAN-based framework aimed at bridging the gap between dance motion and music [4]. Similarly, Li et al. [6] present a transformer-based model featuring a deep cross-modal transformer block with full-attention functionality. Valle-Pérez et al. [7] present a probabilistic autoregressive framework that incorporates both preceding poses and musical information. Huang et al. [13] design a transformer-based system to generate genre-specific dances. Kim et al. [5] propose a conditional GAN framework leveraging transformers to sample latent representations for diverse dance genres. For generating convincing long-term dance sequences, Li et al. [14] devise a two-phase strategy, involving the generation of key poses followed by the prediction of smooth parametric motion curves. Li et al. [15] introduce an architecture with two core elements: a choreographic memory module for encoding and quantizing poses and an actor-critic GPT that sequences these into dances. Gong et al. [16] explore the integration of text and music into dance generation, employing a cross-modal transformer for multimodal feature encoding. Despite these advancements, many existing approaches lack effective mechanisms for flexible editing during motion generation.
Recent works use generative diffusion model to generate diverse motion sequences conditioned on audio features. Tseng et al. [9] propose a diffusion model that offers editing capabilities. Alexanderson et al. [10] apply diffusion models to generate human motion driven by audio, leveraging Conformers for enhanced performance. Li et al. [11] introduce a two-stage diffusion approach that synthesizes long dance movements in a coarse-to-fine manner. Luo et al. [12] introduce a method that improves the spatial-temporal representation of music and dance, achieving a balance between the quality and diversity of generated outputs. Unlike the above approaches, this work concentrates on controllable dance generation by integrating style information within a diffusion framework.
II-B Controllable Human Motion Synthesis
Compared to controllable image/video generation [1, 17], the field of controllable human motion synthesis is less explored. Different from human motion generation [18, 19, 20, 21] that focuses on creating realistic motion sequences, controllable human motion synthesis emphasizes generating motions that adhere to specific constraints or user-defined controls. Karunratanakul et al. [2] propose Guided Motion Diffusion to incorporate spatial constraints into the motion generation process. To enable motion synthesis and editing, Kim et al. [22] propose a diffusion-based model that allows part-based edits at both frame and joint levels without fine-tuning. OmniControl [3] enables flexible, time-specific spatial control over multiple joints in text-conditioned human motion generation using diffusion models. MotionFix [23] generates motion edits as specified by the input text using a conditional diffusion model. DNO [24] optimizes diffusion noise in pre-trained models, enabling efficient motion editing and control without task-specific training. CoMo [25] first decomposes motions into discrete codes and then autoregressively generates sequences of pose codes by leveraging the knowledge priors of large language models. MotionLCM [26] improves real-time controllable motion generation by enhancing runtime efficiency through one-step inference in latent diffusion models. These works focus on text-to-motion, and there is limited research on controllable human motion synthesis conditioned on audio inputs, which this work addresses.
Figure 2: The proposed Style-Guided Motion Diffusion (SGMD). The model is fed with a noisy motion sequence $X_{T}$ of length $N$ in a noising step $T$, along with conditioning music $c$, style prompts $s$ and $T$ itself. The music embedding serves as input for the cross-attention module. During the inference process, SGMD samples noise $X_{T}$ given the conditions $c$ and $s$ to predict the clean sample $\hat{X}$. Subsequently, it diffuses this sample back to $X_{T-1}$, and repeats this iterative process until $t=0$ is reached.
III Method
We introduce Style-Guided Motion Diffusion (SGMD) for controllable dance generation. Enhanced by Spatial-Temporal Masking, the SGMD provides flexibility in generating and editing dance motions. The structure of SGMD is illustrated in Figure 2.
III-A Preliminaries
@@
III-B Style-Guided Motion Diffusion
We extend the Human Motion Diffusion [30] by incorporating style modulation with prompts, resulting in Style-Guided Motion Diffusion (SGMD). In addition to the music condition $c$, this model also includes the style prompts $s$. The training objective is:
where $\hat{x}_{\theta}(\cdot)$ denotes the transformer-based decoder block, which primarily comprises self-attention, feature-wise linear modulation (FiLM) [31], cross-attention, and Style Modulation.
The Style Modulation layer aims to increase the influence of style features on the generated dances. This module, $\mathrm{SM}(\cdot)$, enables the integration of additional style prompts into existing transformer models without affecting the content features. Formally, it is described as:
where $z$ refers to the input, $s$ represents the music style prompts, $r$ is a scaling factor, $\mathrm{FC}(\cdot)$ denotes the fully connected layer, $d_{s}$ is the embedding dimension and $d_{z}$ is the hidden dimension. Unlike adaptive instance normalization [32, 33], which directly manipulates the mean and variance of input and style features to achieve style transfer, the proposed style modulation is more simple and lightweight. The normalized input preserves the content features, while the multiplication incorporates the style features. Since the style modulation module is lightweight and the prompts for a style category are typically pre-extracted, the computational cost introduced by the style prompts is negligible in the diffusion-based framework.
We use classifier-free guidance [34] to train the diffusion model. During training, the condition $c$ and style prompt $s$ are randomly replaced with a low probability. This training strategy enhances the alignment between the generated motions, music conditions, and style prompts. After training, the guided inference is with style prompts is:
$\widetilde{x}(\hat{\sigma}_{t},t,c,s)=w\cdot\hat{x}(\hat{\sigma}_{t},t,c,s)+(1-w)\cdot\hat{x}(\hat{\sigma}_{t},t,\emptyset,s),$ (6)
@@ III-B Style-Guided Motion Diffusion
The overall training loss is the weighted combination defined as follows:
@@
where $\lambda_{j}$, $\lambda_{v}$ and $\lambda_{f}$ are weighted coefficients. The detailed training algorithm is shown in Algorithm 1.
@@
      Add noise on $x$ with $\epsilon$ and predefined parameter $\overline{\alpha}_{t}$: $\sigma_{t}=\sqrt{\overline{\alpha}_{t}}x+(1-\overline{\alpha}_{t})\epsilon$.
@@
      Calculate the losses $L_{j}$, $L_{v}$, and $L_{f}$ in Eq. (1-3).
      Calculate the diffusion loss $L_{d}$ in Eq. (4).
@@
Figure 3: Controllable dance generation with spatial-temporal masking. For the known sequence, we add noise on it to obtain noisy sequence at timestep $t-1$ directly. For unknown sequences, we first use the trained network ${\hat{x}}_{\theta}$ to predict the motion at the timestep 0, and then we add noise on it to obtain noisy sequence at timestep $t-1$ The mask is two-dimensional and allows for control in both the temporal dimension and spatial dimension.
III-C Spatial-Temporal Masking